what is latus rectum?


plzzz explain in deeply.....

Grade:11

1 Answers

vikas askiitian expert
13 years ago

Representation of Latus Rectum of Ellipse

Latus rectum of ellipse

Above diagram shows the graphical representation of Latus Rectum of Ellipse.

Deriving at length of latus rectum of an ellipse

Let LFL’ be a line segment perpendicular to AA’ Meeting the ellipse in L and L’.

Then LL’ is the length of latus rectum of an ellipse .

If LL’ =2l (Length of Latus Rectum of Ellipse) ,

then LF = FL’ = l .

So the coordinate of L are (c,l) .i.e [al,l]

since L(al,l) lies on (x2/a2) + (y2/b2) = 1,

we have [(a2e2/a2)/a2 ] + [l2+b2] = 1 or l2 = b2(1 - e2) = b4/a2[b2(1 - e2 = a2]
Ellipse Latus rectum is

2l=2b2/a